An investor is forming a portfolio by investing $25,000 in stock A which has a beta of 2.40, and $25,000 in stock B which has a beta of 0.60. The return on the market is equal to 8% and treasure bonds have a yield of 3% (rRF). What’s the portfolio beta?

2.40

1.50

1.40

1.80

calculate the required rate of return on the investor’s portfolio

8.0%

10.8%

11.2%

10.5%
Answer #1

HI

Since the weightage of both stock A and stock B is same in portfolio

hence portfolio beta = 0.5*2.4 + 0.5*0.6

= 1.2 + 0.3 = 1.5

Required rate of return can be found out using CAPM

AS per CAPM required return = risk free rate + beta*(market rate - risk free rate)

= 3 + 1.5*(8-3 )

=3+ 1.5* 5

= 10.5%

Thanks